# VLA-Cache: Towards Efficient Vision-Language-Action Model via Adaptive Token Caching in Robotic Manipulation

### v1

@@ Abstract
Vision-Language-Action (VLA) model can process instructions and visual perception to directly generate actions as output in an end-to-end fashion due to its strong multi-modal reasoning capabilities. While the performance of VLA models is promising, their computational cost can be substantial. This raises challenge for applying them on robotics tasks, which requires real-time decision-making to respond quickly to environmental changes. Since robotic control involves sequential decision-making, the visual input often exhibits minimal variation between successive steps. A natural idea is to reuse the computational results of unchanged visual tokens from the last step. Motivated by this idea, we propose VLA-Cache, an efficient vision-language-action model. VLA-Cache incorporates a token-selection mechanism that compares the visual input at each step with the input from the previous step, adaptively identifying visual tokens with minimal changes. The computational results for these unchanged tokens are then reused in subsequent steps via KV-cache, thereby significantly improving the efficiency of the VLA-Cache model. Experimental results on both simulation (e.g., LIBERO benchmark and SIMPLER) and real-world robot valid VLA-Cache can achieve practical acceleration with minimal sacrifice in success rate.
@@ 1 Introduction
Learning a robust and generalizable policy for robotic manipulation through policy learning has long been a challenging problem (Kim et al., 2020), with traditional reinforcement learning approaches (Choi et al., 2024; Bica et al., 2021) often suffering from poor robustness and limited generalization. Recently, the rapid advancement of foundational Vision-Language Models (VLMs) (Awadalla et al., 2023; Liu et al., 2024b) has demonstrated remarkable capabilities in multimodal understanding and generalization. Leveraging large-scale real-world robotic datasets (O’Neill et al., 2024; Fang et al., 2024), pioneering works (Zitkovich et al., 2023; Octo Model Team et al., 2024; Niu et al., 2024; Kim et al., 2024) have introduced Vision-Language-Action (VLA) models, which integrate vision and language modalities to directly generate robotic actions in an end-to-end manner. This emerging paradigm holds great promise for enhancing the adaptability and generalization of robotic control systems, but leaves a large computational demand.
@@
Figure 1: During the inference of the VLA model, static tokens of the input image remain largely consistent across steps. This consistency allows for caching the computations of these tokens from the previous step and improving the efficiency of inference.
To mitigate the extensive cost of VLA models, existing works often adopt generic acceleration techniques, such as model lightweighting (Wen et al., 2024), quantization (Park et al., 2024), and early exit (Yue et al., 2024). However, these approaches fail to account for the unique characteristics of VLA models, leading to suboptimal efficiency gains and limited performance improvements.
In this paper, we argue that robotic manipulation, characterized by sequential action-scene interactions, inherently prioritizes visual inputs related to the robotic arm and the target object, while the background remains largely static and less critical. As illustrated in Figure 1, the static tokens of the input image remain largely consistent across steps. In this context, the background tokens, which contribute significantly to computational overhead and repeat across sequential steps in the VLA process, are often redundant.
As a result, we propose VLA-Cache, a novel approach that identifies and reuses static tokens across sequential VLA steps. Specifically, static tokens in each step refer to visual tokens that exhibit minor changes compared to the previous step. Instead of recomputing these tokens, VLA-Cache directly substitutes them with their cached representations from the preceding step, significantly reducing redundant computations and improving inference efficiency. Nevertheless, we observe that model precision is highly sensitive to task-relevant tokens (i.e., those related to the robotic arm and the target object) in the ablation study, where even a slight positional shift can lead to a significant performance drop. To address this, we introduce a fine-grained selection scheme that filters out task-relevant tokens from the static token set, ensuring that these critical tokens continue to undergo full computational interactions, thereby preserving model accuracy while maintaining efficiency.
Additionally, based on the observation that attention distributions within the VLA decoder vary significantly across different layers, we propose a layer-adaptive strategy that adjusts the fraction of reused tokens based on each layer’s attention concentration. We evaluate VLA-Cache on a diverse set of robotic manipulation tasks in the LIBERO benchmark, where it demonstrates over 1.7× acceleration compared to the baseline, with only slight drop on the success rate. Furthermore, VLA-Cache is deployed on the Kinova Jaco robot arm, showing that it can achieve practical acceleration in real-world scenarios.
@@ 2 Related Work
Vision-Language-Action Models. Large-scale vision-language models (VLMs) have significantly advanced multimodal learning by integrating image understanding and language reasoning (Liu et al., 2024b; Bai et al., 2023). Extending these capabilities, Vision-Language-Action (VLA) models (Zitkovich et al., 2023; Li et al., 2023) incorporate an action modality, enabling end-to-end visuomotor control. These models typically adopt large VLM backbones (Touvron et al., 2023) and fine-tune them on robot data (O’Neill et al., 2024), with approaches varying from discretizing actions as language-like tokens (Kim et al., 2024; Chi et al., 2023) to incorporating specialized diffusion policy heads (Black et al., 2024). Despite their effectiveness in tasks like object retrieval and assembly (Huang et al., 2024; Zhao et al., 2023), VLA models demand substantial computation, making real-time deployment challenging, particularly in resource-constrained environments.
Acceleration for Vision-Language Models. Acceleration techniques for VLMs, such as quantization (Xie et al., 2024) and pruning (Lin et al., 2024a), have shown success in vision-language tasks but often overlook the real-time demands of action generation. Token-level methods like FastV (Chen et al., 2025) and SparseVLM (Zhang et al., 2024) optimize inference by pruning or merging redundant tokens, but their effectiveness in VLA tasks, which require short and highly specialized action sequences, remains unclear. Some works address efficiency by modifying VLA architectures, such as DeeR-VLA (Yue et al., 2024), which dynamically adjusts inference depth, and QAIL (Park et al., 2024), which integrates quantization-aware training. Others, like RoboMamba (Liu et al., 2024c) and TinyVLA (Wen et al., 2024), focus on replacing traditional attention mechanisms or training compact models from scratch, often requiring re-training and additional data collection. In contrast, VLA-Cache offers a training-free acceleration approach by selectively caching static tokens and recomputing only dynamic or task-relevant ones, preserving crucial temporal cues while reducing inference latency and improving real-time robotic control efficiency.
3 Preliminary
3.1 Vision-Language-Action Models
Vision-Language-Action (VLA) models extend large-scale VLMs to generate robotic actions based on multimodal inputs, including visual observations and task instructions. These models are typically trained on large-scale datasets, such as Open X Embodiment (O’Neill et al., 2024), which contains over a million real-world robot trajectories. The training pipeline involves a pre-trained vision encoder (Liu et al., 2021; Radford et al., 2021) extracting image features, which are projected into the input space of a Large Language Model (LLM). The LLM, fine-tuned alongside a projector module, encodes these inputs to predict actions.
During inference, VLA models process sequential frames to iteratively guide the robot, but the full Transformer-based LLM inference at each step results in substantial computational overhead. The quadratic complexity of self-attention, combined with large key-value (KV) caches (Ge et al., 2023), leads to high memory consumption and slow inference speeds, posing challenges for real-time deployment. While model compression techniques such as quantization (Lin et al., 2024b) and pruning (Park et al., 2024; Yue et al., 2024) have been explored to reduce computational cost, they typically require re-training and may degrade performance, limiting their practicality in real-time applications. To mitigate these inefficiencies, our approach leverages token caching mechanisms to accelerate inference in a training-free manner, reducing computational cost while preserving inference accuracy.
3.2 Key-Value Caching in Transformer Models
Key-Value (KV) caching is a well-established optimization in autoregressive Transformers (Vaswani, 2017; Floridi & Chiriatti, 2020), allowing previously computed key ($\mathbf{K}$) and value ($\mathbf{V}$) representations to be stored and reused in subsequent timesteps. Given input tokens $\mathbf{X}$, self-attention first projects them into queries, keys, and values:
$\mathbf{Q}=\mathbf{X}W_{Q},\quad\mathbf{K}=\mathbf{X}W_{K},\quad\mathbf{V}=\mathbf{X}W_{V}.$ (1)
The attention output is computed as:
@@ 3.2 Key-Value Caching in Transformer Models
In standard inference, new keys and values are appended to the cache:
$\mathbf{K}_{t}=\mathrm{Concat}(\mathbf{K}_{t-1},\mathbf{K}_{\text{new}}),\quad\mathbf{V}_{t}=\mathrm{Concat}(\mathbf{V}_{t-1},\mathbf{V}_{\text{new}}).$ (3)
While this improves efficiency by avoiding redundant recomputation, it does not reduce memory or computational costs, as all tokens are still processed.
VLA-Cache extends this approach by introducing dynamic token reuse. Instead of appending all new tokens, it identifies static tokens across frames and skips their recomputation:
$\mathbf{K}_{t}(i)=\begin{cases}\mathbf{K}_{t-1}(i),&i\in\mathcal{P}_{\text{reuse}},\\ W_{K}\mathbf{H}_{t}(i),&\text{otherwise}.\end{cases}$ (4)
$\mathbf{V}_{t}(i)=\begin{cases}\mathbf{V}_{t-1}(i),&i\in\mathcal{P}_{\text{reuse}},\\ W_{V}\mathbf{H}_{t}(i),&\text{otherwise}.\end{cases}$ (5)
where $\mathcal{P}_{\text{reuse}}$ represents the subset of tokens that remain visually and semantically unchanged between frames. By selectively updating only task-relevant tokens, VLA-Cache significantly reduces redundant computation while preserving high-fidelity action predictions.
This caching mechanism integrates seamlessly into the Transformer decoding process, requiring no architectural modifications or additional training. The following sections detail the implementation of VLA-Cache and its effectiveness in improving inference efficiency for VLA models.
Figure 2: In this paper, we propose VLA-Cache, which includes two procedures. (a) Dynamic Token Selection: First, a static token set is identified, followed by filtering out task-relevant tokens to ensure critical information undergoes full computation. Tokens that remain in the set and overlap with those from the previous step will be retrieved from cache. (b) Adaptive Token Caching: The fraction of reused tokens is dynamically adjusted at each layer based on its attention distribution.
4 Methodology
In robotic action prediction, consecutive image frames often show only minor changes in the manipulator or target objects, while the background remains mostly unchanged. This motivates us to reuse some of unchanged tokens in the last step to accelerate the inference at current step. However, the choice of reusing tokens should be made very carefully, as even small changes to the manipulator or target objects can be pivotal for accurate action prediction. Treating such tokens as static could severely degrade performance. Consequently, the core challenge in accelerating VLA inference via token caching lies in accurately distinguishing genuinely static regions from those that are dynamic or closely tied to the task. By automatically detecting static tokens across consecutive frames and discarding those related to the manipulator or the immediate task, one can realize substantial computational savings while still preserving high-quality action predictions.
4.1 Dynamic Token Selection
Static Token Selection. To enhance inference efficiency, we identify visually static tokens across consecutive frames and reuse their associated Key-Value (KV) representations. This process involves three steps: partitioning the input image into patches, computing patch-wise similarity, and selecting the most static patches for reuse.
@@ 4.1 Dynamic Token Selection
Given an image $I$ of size $H\times W$, we divide it into $N\times N$ non-overlapping patches of size $p\times p$, yielding $N^{2}$ patches $\mathcal{P}=\{\mathbf{P}_{i,j}\}$. Each patch $\mathbf{P}_{i,j}$ is directly represented by its raw pixel values, without additional feature encoding.
For two consecutive frames $I_{t}$ and $I_{t-1}$, we calculate the cosine similarity between their corresponding patches:
$\text{Sim}\big{(}\mathbf{P}_{t}^{i,j},\mathbf{P}_{t-1}^{i,j}\big{)}=\frac{\mathbf{P}_{t}^{i,j}\cdot\mathbf{P}_{t-1}^{i,j}}{\|\mathbf{P}_{t}^{i,j}\|_{2}\,\|\mathbf{P}_{t-1}^{i,j}\|_{2}}.$ (6)
@@
A patch is considered static if its similarity score exceeds a predefined threshold $\tau$. To further refine the selection, we apply a Top-$k$ filter:
$\mathcal{P}_{\mathrm{top\text{-}k}}=\mathrm{Top}\text{-}k\Big{(}\{\mathbf{P}_{t}^{i,j}\mid\mathrm{Sim}(\mathbf{P}_{t}^{i,j},\mathbf{P}_{t-1}^{i,j})\geq\tau\}\Big{)}.$ (7)
@@
Thus, patches with minimal visual variation are identified for reuse, while those displaying notable differences are recomputed. This approach accurately selects truly static tokens across consecutive frames, significantly reducing redundant computations and accelerating inference without compromising overall performance.
Evicting Task-Relevant Tokens. While many tokens remain visually static across consecutive frames, certain tokens that are highly relevant to the task (e.g., the gripper or the target object) cannot be directly reused. In our experiments, we observed that as the robotic arm approaches the target object, small but critical changes in the target region often occur. These changes are primarily caused by factors such as lighting variations or camera noise. Recomputing tokens around the target region before the gripper reaches it significantly improves task accuracy. To address this issue, we leverage the attention mechanism of the VLA model to identify task-relevant tokens in the next frame that require recalculation.
Specifically, we calculate the text-to-vision attention scores from the cross-attention module of the VLA model to determine the task relevance of each vision token. For each layer $l$, the attention weights are represented as $\mathbf{A}^{l}\in\mathbb{R}^{N_{\text{heads}}\times N_{\text{tokens}}\times N_{\text{tokens}}}$. From these weights, we extract the attention mapping from text tokens to vision tokens:

### v2

# VLA-Cache: Efficient Vision-Language-Action Manipulation via Adaptive Token Caching

@@ Abstract
Vision-Language-Action (VLA) models have demonstrated strong multi-modal reasoning capabilities, enabling direct action generation from visual perception and language instructions in an end-to-end manner. However, their substantial computational cost poses a challenge for real-time robotic control, where rapid decision-making is essential. This paper introduces VLA-Cache, a training-free inference acceleration method that reduces computational overhead by adaptively caching and reusing static visual tokens across frames. Exploiting the temporal continuity in robotic manipulation, VLA-Cache identifies minimally changed tokens between adjacent frames and reuses their cached key-value representations, thereby circumventing redundant computations. Additionally, to maintain action precision, VLA-Cache selectively re-computes task-relevant tokens that are environmentally sensitive, ensuring the fidelity of critical visual information. To further optimize efficiency, we introduce a layer adaptive token reusing strategy that dynamically adjusts the reuse ratio based on attention concentration across decoder layers, prioritizing critical tokens for recomputation. Extensive experiments on two simulation platforms (LIBERO and SIMPLER) and a real-world robotic system demonstrate that VLA-Cache achieves up to 1.7× speedup in CUDA latency and a 15% increase in control frequency, with negligible loss on task success rate. The code and videos can be found at our project page: https://vla-cache.github.io.
@@ 1 Introduction
Learning a robust and generalizable policy for robotic manipulation through policy learning has long been a challenging problem [1], with traditional reinforcement learning approaches [2, 3] often suffering from poor robustness and limited generalization. Recently, the rapid advancement of foundational Vision-Language Models (VLMs) [4, 5] has demonstrated remarkable capabilities in multimodal understanding and generalization. Leveraging large-scale real-world robotic datasets [6, 7], pioneering works [8, 9, 10, 11] have introduced Vision-Language-Action (VLA) models, which integrate vision and language modalities to directly generate robotic actions in an end-to-end manner. This emerging paradigm holds great promise for enhancing the adaptability and generalization of robotic control systems, but leaves a large computational demand.
To mitigate the extensive cost of VLA models, existing works often adopt generic acceleration techniques, such as model lightweighting [12], quantization [13], and early-exit [14]. While effective to some extent, these methods often require architectural modifications or retraining, and more importantly, they lack task-specific design tailored to the intrinsic characteristics of VLA tasks. As a result, they struggle to achieve an optimal balance between inference speed and action accuracy. In this paper, we step into the nature of VLA robotic manipulation that is intrinsically different from the VLM models. Specifically, VLA tasks involve sequentially processing a stream of temporally adjacent visual observations, where the environment often exhibits high spatial redundancy across time. As shown in Figure 1, large portions of the visual scene, especially background regions, remain static and semantically irrelevant to action decisions, yet are processed repeatedly at each time step. These static tokens contribute significantly to computational overhead while providing limited utility for downstream control. This motivates our proposed token caching mechanism, which explicitly exploits temporal redundancy in visual inputs to reduce redundant computation without compromising decision quality.
@@
Figure 1: During the inference of the VLA model, static tokens of the input image remain largely consistent across steps. This consistency allows for caching the computations of these tokens from the previous step.
To address the inefficiency introduced by repeatedly processing static visual information, we present VLA-Cache, a training-free inference acceleration method that exploits temporal continuity in robotic perception. Rather than recomputing all vision tokens at every timestep, VLA-Cache identifies tokens that exhibit minimal change between adjacent frames and reuses their cached key-value (KV) representations to bypass redundant computation. However, we observe that not all visually static tokens can be safely reused. Some tokens, such as those near the gripper or target object, may appear visually unchanged but remain semantically active and crucial for accurate action generation. Naively reusing all static tokens results in a significant performance drop as shown in Table  1. To mitigate this, VLA-Cache incorporates a lightweight filtering mechanism based on decoder attention scores to exclude task-relevant tokens from reuse, ensuring that semantically critical regions are always recomputed with up-to-date features. Moreover, we observe that attention patterns vary across decoder layers, with deeper layers exhibiting more concentrated focus. To further optimize reuse, VLA-Cache employs a layer-adaptive caching strategy that dynamically adjusts the reuse ratio per layer based on attention entropy, prioritizing precise updates in sensitive regions. These two mechanisms together enable substantial reduction in decoding overhead, especially in large-scale language decoders (e.g., LLaMA [15], Gemma [16]), which typically dominate the compute cost in VLA systems.
[…]
@@ 2 Related Work
Vision-Language-Action Models. Large-scale vision-language models (VLMs) have significantly advanced multimodal learning by integrating image understanding and language reasoning [5, 21]. Extending these capabilities, VLA models [8, 22] incorporate an action modality, enabling end-to-end visuomotor control. These models typically adopt large VLM backbones [15] and fine-tune them on robot data [6], with approaches varying from discretizing actions as language-like tokens [11, 23] to incorporating specialized diffusion policy heads [24]. Despite their effectiveness in tasks like object retrieval and assembly [25, 26], VLA models demand substantial computation, making real-time deployment challenging, particularly in resource-constrained environments.
Acceleration for Vision-Language Models. Inference acceleration has been extensively explored in vision-language models (VLMs) through quantization [27], pruning [28], and token-level techniques such as FastV [29], SparseVLM [30], ToMe [31], PuMer [32], and MADTP [33]. These intra-frame strategies reduce redundancy within a single image but disregard the temporal and spatial structure essential for robotic tasks under closed-loop control. In the VLA domain, efficiency has been addressed through architectural modifications (e.g., RoboMamba [34], TinyVLA [12]), quantization-aware training (QAIL [13]), and dynamic depth control (DeeR-VLA [14]). While effective, these methods require re-training and lack generalizability. Recent high-frequency frameworks such as $\pi_{0}$-FAST [35], HiRT [36], and OpenVLA-OFT [20] achieve higher control frequency via action chunking or asynchronous decoding. However, they continue to suffer from the language model decoding bottleneck, which dominates inference time. VLA-Cache addresses this gap by introducing a cross-frame token reuse strategy that accelerates inference without modifying the model or requiring additional training. Furthermore, it complements existing high-frequency VLA architectures by directly accelerating the language decoder bottleneck, offering a lightweight, plug-and-play solution for real-time robotic inference.
3 Methodology
In robotic action prediction, most visual tokens remain static across frames except for key regions like the manipulator or target object. While this temporal redundancy enables token reuse, reusing all static tokens can harm accuracy when task-relevant regions subtly change. To address this, we propose a method that identifies visually static tokens and filters out semantically important ones based on attention scores from the VLA decoder. By avoiding redundant computation of unchanged static tokens between adjacent frames, our approach directly alleviates the computation bottleneck of language decoder in VLA models while preserving the accuracy of action prediction.
3.1 KV Cache for VLA Token Reusing
Key-Value (KV) caching is a widely adopted technique in large-scale autoregressive models to reduce both computation and memory footprint during decoding. Initially proposed in the Transformer architecture [37], KV caching enables the model to reuse previously computed key ($\mathbf{K}$) and value ($\mathbf{V}$) vectors for each token, thereby avoiding redundant computation across decoding steps. Concretely, given a sequence of input tokens $\mathbf{X}$, the self-attention mechanism computes:
$\mathbf{Q}=\mathbf{X}W_{Q},\quad\mathbf{K}=\mathbf{X}W_{K},\quad\mathbf{V}=\mathbf{X}W_{V},$ (1)
@@ 3.1 KV Cache for VLA Token Reusing
During decoding, each new token’s $\mathbf{k}_{\text{new}}$ and $\mathbf{v}_{\text{new}}$ are appended to the $i$-th cache:
$\mathbf{K}_{i}=\mathrm{Concat}(\mathbf{K}_{i-1},\mathbf{k}_{\text{new}}),\quad\mathbf{V}_{i}=\mathrm{Concat}(\mathbf{V}_{i-1},\mathbf{v}_{\text{new}}).$ (3)
While KV caching is effective for language decoding within a single query in vision-language models, this technique does not address redundancy in the visual stream, especially in Vision-Language-Action (VLA) models. In robotic manipulation, consecutive visual inputs often share large overlapping content, yet VLA models typically discard visual encodings after each step and recompute them from scratch. This is both wasteful and suboptimal for real-time control.
This inefficiency motivates a key question: can we selectively reuse static visual tokens across time with temporal KV Caching in VLA models? This idea forms the basis of our proposed VLA-Cache. In the following sections, we introduce its core mechanisms: static token selection, task-relevance filtering, and layer-adaptive reuse to accelerate VLA inference while preserving action accuracy.
Figure 2: VLA-Cache accelerates the VLA’s language decoding process across timesteps via the following two steps: (a) Dynamic Token Selection reuses static tokens across frames while preserving task-relevant ones; (b) Adaptive Token Caching dynamically adjusts reuse ratios per decoder layer based on attention patterns.
3.2 Temporal Redundancy in Robotic Perception
In closed-loop robotic manipulation, consecutive visual frames often share large portions of static content. As illustrated in Figure 2, background regions and stationary objects typically exhibit negligible changes between adjacent frames. However, most existing Vision-Language-Action (VLA) models discard visual representations after each timestep and recompute all visual tokens from scratch, resulting in substantial redundancy and increased inference latency.
To address this inefficiency, we propose selectively reusing visual tokens that remain static across timesteps. Specifically, we identify and cache the representations of image regions with minimal visual change, allowing their Key-Value (KV) representations to be reused in the next frame. This strategy significantly reduces redundant computation in the VLA visual stream while preserving model performance.
@@ 3.2 Temporal Redundancy in Robotic Perception
Static Token Selection. Given an image $I\in\mathbb{R}^{H\times W\times 3}$, we divide it into $N\times N$ non-overlapping patches of size $p\times p$, yielding a set of raw pixel patches $\mathcal{P}_{t}=\{\mathbf{P}_{t}^{i,j}\}$. For each patch $\mathbf{P}_{t}^{i,j}$ in the current frame and its corresponding patch $\mathbf{P}_{t-1}^{i,j}$ in the previous frame, we compute cosine similarity:
$\text{Sim}\big(\mathbf{P}_{t}^{i,j},\mathbf{P}_{t-1}^{i,j}\big)=\frac{\mathbf{P}_{t}^{i,j}\cdot\mathbf{P}_{t-1}^{i,j}}{\|\mathbf{P}_{t}^{i,j}\|_{2}\cdot\|\mathbf{P}_{t-1}^{i,j}\|_{2}}.$ (4)
@@
A patch is considered visually static if its similarity exceeds a threshold $\tau$. We further apply a Top-$k$ filter to retain the most stable tokens:
$\mathcal{P}_{\mathrm{static}}=\mathrm{Top}\text{-}k\Big(\{\mathbf{P}_{t}^{i,j}\mid\mathrm{Sim}(\mathbf{P}_{t}^{i,j},\mathbf{P}_{t-1}^{i,j})\geq\tau\}\Big).$ (5)
@@
This simple yet effective approach accurately selects truly static tokens across consecutive frames, significantly reducing redundant computations and accelerating inference without compromising overall performance.
Table 1: Comparison of VLA-Cache’s core token selection strategies on OpenVLA using the LIBERO Spatial benchmark.
Method SR (%) $\uparrow$ Latency (ms) $\downarrow$
OpenVLA 84.4 51.56
+ Static Token 74.2 31.03
+ Evict Task-Relevant 82.6 31.03
+ Layer Adaptive 83.8 32.22
3.3 Retaining Task-Relevant Information
While visual similarity offers a practical signal for reusing static regions, not all visually static tokens are safe to reuse. In robotic control tasks, certain regions, such as the gripper or target object, though visually unchanged, are semantically dynamic and critical for precise action generation. Directly reusing all visually static tokens, without considering their semantic role, can lead to serious performance degradation. As shown in Table 1, naive static token reuse lowers the success rate from the OpenVLA baseline of 84.4% to just 74.2%.
This degradation occurs because task-relevant tokens, though minimally changed at the pixel level, exhibit greater feature sensitivity to subtle environmental changes. Unlike vision-language models (VLMs), VLA models must track object states and interactions over time, making them more dependent on accurate visual encoding. Therefore, to ensure alignment with the latest environment state, task-relevant regions must be recomputed each step.
Evicting Task-Relevant Tokens. To avoid reusing semantically critical but visually static tokens, we propose a lightweight filtering mechanism using cross-attention scores from the language decoder. For each decoder layer $l$, we extract the text-to-vision attention matrix $\mathbf{A}^{l}_{\text{vis-text}}$ from the full attention tensor $\mathbf{A}^{l}\in\mathbb{R}^{N_{\text{heads}}\times N_{\text{tokens}}\times N_{\text{tokens}}}$ as: